OpenStudy (anonymous):

Find a number that is between 2/7 and 4/11 .

OpenStudy (anonymous):

Make denominators equal first..

OpenStudy (anonymous):

ok so what is it

OpenStudy (anonymous):

@naranjja you were writing something. Carry on buddy..

OpenStudy (anonymous):

So @Hero ?

hero (hero):

\[\frac{2}{7} and \frac{4}{11}\] After cross multiplying notice that 22 < 28 So \[\frac{2}{7} < \frac{4}{11}\] Now \(\dfrac{3}{7} > \dfrac{2}{7}\) So test if \(\dfrac{3}{7} < \dfrac{4}{11}\) If so, then you have found a fraction between \[\frac{2}{7} and \frac{4}{11}\]

OpenStudy (anonymous):

My browser crashed. Sorry, I'll start over again.

hero (hero):

Okay, so that didn't work. But now you can test \(\dfrac{3}{11}\) We know that \(\dfrac{3}{11} < \dfrac{4}{11}\) So test to see if \(\dfrac{3}{11} > \dfrac{2}{7}\) That doesn't work either. But you can try adding both fractions together then dividing by 2 to find a fraction between both: \[\left(\frac{2}{7} + \frac{4}{11}\right) \div 2\]

OpenStudy (anonymous):

Establishing the common denominator 77 (multiplying the first by 11 and the second by 7), \[\frac{ 2 }{ 7 }=\frac{ 22 }{ 77 }\]\[\frac{ 4 }{ 11 }=\frac{ 28 }{ 77 }\] So say you choose, \[\frac{ 24 }{ 77 }\] That fraction is irreducible, and it is approximately 0.312

hero (hero):

That's probaby the easiest way to do it.

OpenStudy (anonymous):

\[\frac{ 22 }{ 77 }\approx0.28\]\[\frac{ 28 }{ 77 }\approx0.36\] You can choose 0.3 for example, which is 1/3

OpenStudy (anonymous):

so is it 22/77

hero (hero):

@naranjja, there is no need to reduce further. If you have two fractions \(\dfrac{a}{c}\) and \(\dfrac{b}{c}\) and a < b, then \(\dfrac{a}{c}\) < \(\dfrac{b}{c}\)

hero (hero):

and if you find a third fraction \(\dfrac{d}{c}\) such that a < d < c, then \(\dfrac{a}{c} < \dfrac{d}{c} < \dfrac{b}{c}\)

OpenStudy (anonymous):

soooo

OpenStudy (anonymous):

is it 22/77?

hero (hero):

No, it isn't. I guess you haven't understood anything me or @naranjja has written.

OpenStudy (anonymous):

It can be any number between 22/77 and 28/77. Which is the same as saying between 0.28 and 0.36. I think 0.333 is a good number to choose because it is 1/3.

OpenStudy (anonymous):

oh ok and yes I did just making sure @Hero

OpenStudy (anonymous):